OpenStudy (anonymous):

Which points are 56 units from the point located at 13 on a number line? A. –69 and 69 B. –69 and 43 C. –43 and 69 D. 43 and 69

OpenStudy (anonymous):

@Pawtpie

OpenStudy (anonymous):

So if you can imagine the number 13 on a number line, we need the number located at 56 units in the positive direction, and 56 units in the negative direction.

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

What is: 13 + 56 and 13 - 56

OpenStudy (anonymous):

13+56=69 13-56=-43

OpenStudy (anonymous):

Which is answer choice? :)

OpenStudy (anonymous):

C. -46 and 69

OpenStudy (anonymous):

Now I may say, you are correct :).

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

You said -46 though, I'm sure you meant -43?

OpenStudy (anonymous):

no I meant -46

OpenStudy (anonymous):

:}

OpenStudy (anonymous):

13 - 56 should = -43

OpenStudy (anonymous):

u r right

OpenStudy (anonymous):

With that, your answer choice C seems correct

OpenStudy (anonymous):

sorry that I confused u

OpenStudy (anonymous):

Oh no, nothing to be sorry for :)

OpenStudy (anonymous):

if I could I would give u 5 medals 4 this one